OpenStudy (anonymous):

Solve each equation for the indicated variable: F=9/5c+32 Solve for C

OpenStudy (asnaseer):

is this your question?\[F=\frac{9}{5}C+32\]

OpenStudy (anonymous):

the order of operations says you must subtract the 32 from both sides first, if you want to solve for C: F = (9/5)C + 32 F - 32 = (9/5)C (5/9)(F - 32) = C